# Understanding Before Recommendation: Semantic Aspect-Aware Review Exploitation via Large Language Models

### v1

@@ 4.1. Experimental setup
Amazon Review Datasets  [^2] http://jmcauley.ucsd.edu/data/amazon stems from the extensive product review system of Amazon (Liu et al., 2019). It includes user-product interactions, reviews, and metadata. In our experiments, only the interaction information is utilized. Three product categories from the 5-core version of this dataset are selected for evaluation: Office, Baby, and Clothing.
For each dataset, we randomly split the interaction of each user into training and testing sets with a ratio of 80:20. From the training set, we randomly select 10% of interactions to form the validation set for tuning hyper-parameters. The pairwise learning strategy is adopted for parameter optimization. Specifically, the observed user-item interactions in the training set were treated as positive instances. For each positive instance, we pair it with a random sampling negative instance that the user has not interacted with yet.
4.1.2. Compared Baselines.
To demonstrate the effectiveness of ClusterGCF, we compared it with several baseline models.
  • NeuMF (He et al., 2017): It is a state-of-the-art neural collaborative filtering method that introduces multiple hidden layers to capture the non-linear interactions between their features.
  • GCMC (van den Berg et al., 2018): This method leverages the GCN technique (Kipf and Welling, 2017) on the user-item bipartite graph and employs one convolutional layer to exploit direct connections between users and items.
  • LightGCN (He et al., 2020): This model can be considered a simplified version of NGCF (Wang et al., 2019b), where certain components like feature transformation and nonlinear activation module are removed.
  • NCL (Lin et al., 2022): This is a model-agnostic contrastive learning framework. It seeks to capture the correlation between a node and its prototype to improve the neural graph collaborative filtering.
  • LATTICE (Fan et al., 2022): This model uses side information to construct an item-item graph and perform graph convolution operations to aggregate information from high-order neighbors.
  • MMGCN (Wei et al., 2019): The model learns user preferences and item characteristics by propagating modality embeddings on the user-item graph.
  • GRCN (Wei et al., 2020): This GCN-based multimedia recommendation model adjusts the structure of the interaction graph adaptively based on the status of model training for discovering and pruning potential false-positive edges.
  • BM3 (Zhou et al., 2023): This self-supervised framework uses a latent representation dropout mechanism to generate the target view of a user or an item. BM3 is trained without negative samples by jointly optimizing three multi-modal objectives.
For fair comparisons, all methods are optimized using the same pairwise learning strategy. We put great efforts into tuning these methods based on the validation dataset and reported their best performance.
4.1.3. Evaluation Metrics
Two widely-used evaluation metrics are adopted for top-$K$ recommendation: Recall and Normalized Discounted Cumulative Gain (NDCG) (He et al., 2015). Recommendation accuracy is calculated for each metric based on the top 10 and 20 results. Note that the reported results are the average values across all testing users.
4.1.4. Parameter Settings
The PyTorch framework  [^3] https://pytorch.org. is adopted to implement the proposed model. Vicuna-13B  [^4] https://huggingface.co/lmsys/vicuna-13b-v1.5 is used to analyze the reviews. In our experiments, all hyperparameters are carefully tuned. The user and item embedding size of each aspect is set to 64 across all datasets. The mini-batch size is fixed to 1024. The learning rate for the optimizer is searched from $\{1e^{-5},1e^{-4},\cdots,1e^{1}\}$, and the model weight decay is searched in the range $\{1e^{-5},1e^{-4},\cdots,1e^{-1}\}$. Besides, the early stopping strategy is adopted. Specifically, the training process will stop if recall@10 does not increase for 20 successive.
@@ 4.1.4. Parameter Settings
Table 3. The extracted semantic-aware aspects and the number of their associated interactions over three datasets.
@@
4.2. Performance Comparison
In this section, we compared the performance of our proposed method with all the adopted competitors. The experimental results (in terms of Recall@10, Recall@20, NDCG@10, and NDCG@20) of our method and all compared methods over three test datasets are reported in Table 2. Note that the best and second-best results are highlighted in bold. From the results, we have the following observations.
We first focus on the performance of methods in the first block, which only uses the user-item interaction information. NeuMF is a deep learning-based method, and it shows relatively poor performance. GCMC performs better than NeuMF, highlighting the benefits of leveraging graph structure. However, GCMC underperforms LightGCN because it only utilizes first-order neighbor nodes. This showcases the strength of GCN and the importance of incorporating high-order information into representation learning. Moreover, DGCF performs better than LightGCN by leveraging the disentangled representation technique to model users’ diverse intents. Both IMP-GCN and NCL consistently outperform all the above GCN-based methods over all the datasets. This is mainly due to their ability to filter out noisy information from neighboring nodes effectively. In particular, IMP-GCN’s performance superiority over LightGCN and DGCF demonstrates the importance of distinguishing nodes in high-order neighbors in the graph convolution operation.
The second block is the GCN-based methods utilizing side information. MMGCN is the first method combining side information with GCN, but it was found to be less effective than GRCN. This is attributed to the GRCN adaptively adjusting the false-positive edges of the interaction graph to discharge the ability of GCN. LATTICE outperforms GRCN on Baby and Clothing by constructing item-item graphs based on item similarity in the side information. BM3 consistently performs better than all general CF baselines over Baby and Clothing. It should be credited to the joint optimization of multiple objectives to learn the representations of users and items.
SAGCN consistently yields the best performance across all datasets. In particular, the improvements of SAGCN over the strongest competitors $w.r.t$ NDCG@10 are 15.77%, 15.89%, and 21.77% in Office, Baby, and Clothing, respectively. We credit this to the combined effects of the following three aspects. Firstly, we accurately extract structured semantic aspect-aware reviews using LLM to understand user behaviors at a finer level, which also avoids the noise in raw reviews. Secondly, the SAGCN performs graph convolution on various semantic-aspect-based graphs, which can alleviate the over-smoothing problem (Liu et al., 2021) and avoid negative information from unreliable neighbor nodes (Wei et al., 2020). Thirdly, we extract fine-grained user behaviors for each semantic aspect to bridge the modality gap between interaction and textual features.
4.3. Effect of Layer Numbers
(a) Recall on Office
(b) Recall on Clothing
Figure 4. Performance Comparison between SAGCN and competitors at different layers on Office and Clothing.
4.4. Semantic Aspects Analysis
@@ 4.4.1. Semantic Aspect-Aware Interactions Analysis
In our experiment, we only selected eight aspects per dataset due to sparse interactions in other aspects. Table 3 reported the number of interactions for different semantic aspects across three datasets.
Based on our observations of various recommendation scenarios, it is evident that users have distinct concerns regarding different aspects. For instance, in reviews of office products, convenience is a common concern, while fit is only mentioned in clothing reviews. To rank the semantic aspects in each recommendation scenario, we analyzed their interaction numbers. This helps us identify that the importance of different semantic aspects varies depending on the number of interactions. For example, in reviews of Office products, Durability was mentioned more frequently than Design. On the other hand, in reviews of Baby, Design was mentioned more frequently than Durability.
@@ 4.4.2. Effect of Aspect Numbers
(a) Recall on Office
(b) Recall on Clothing
To investigate the effect of semantic aspect numbers, we conducted experiments on Office and Clothing. Note that all aspects were ranked according to the number of interactions, we selected a number of aspects based on their ranking. Our findings, shown in Fig. 5, suggest that increasing the semantic aspect numbers can improve the performance of SAGCN. In other words, SAGCN achieves its peak performance when leveraging all aspects, indicating that it can benefit from various aspects. Furthermore, we found that SAGCN with eight aspects showed a greater improvement in performance than it with six aspects. This could be due to the fact that the added two aspects might provide more supplementary information to the other six. When combined, the full set of features gives the model a richer understanding of user behaviors.
@@ 4.5. Interpretability Analysis
Figure 6. Performance comparison of SAGCN on different semantic aspects on Office and Clothing. Notice that the values are reported by percentage with ’%’ omitted.
@@
Figure 7. The preference score of user $u3547$ for item $i836$ and $i1322$ from different semantic aspects on Office. Notice that the values are reported by percentage with ’%’ omitted.
In this section, we examined the interpretability of our model from two different perspectives.
@@
First, our model can interpret the contribution of each semantic aspect in a recommendation scenario. To examine the contribution of different aspects, we predicted the users’ preferences using individual aspect embeddings of the users and items. Fig. 6 shows the following interesting observations: (1) Different semantic aspect features have varying contributions in recommendation. For instance, Quality has a greater impact than Design for Office. (2) The number of semantic-aware interactions is not the main contribution factor. Surprisingly, even though the number of interactions for Price is the least, it still has the highest contribution to our model. This could be because Price introduces more valuable interactions than the other aspects.
@@
Second, we can further provide a detailed explanation by providing the scores of various semantic aspects for each user-item pair. In Fig. 7, we computed the preference score of user $u3547$ for item $i836$ and $i1322$ from different semantic aspects on Office. From the result, we can conclude that the user would be interested in these two items mainly due to their Convenience and Quality. Additionally, the Comfort aspect of $i1322$ might also attract the user’s attention.
5. Conclusion

### v2

@@ 4.2. Performance Comparison With Baseline Methods
In this section, we compared the performance of our proposed SAGCN with all the adopted competitors (conventional recommendation models). The experimental results of our method and all compared methods over four test datasets in terms of Recall@10, Recall@20, NDCG@10, and NDCG@20 are reported in Table 2. Note that the best and second-best results are highlighted in bold. From the results, we have the following observations.
We first focus on the performance of methods in the first block, which only uses the user-item interaction information. NeuMF is a deep learning-based method, and it shows relatively poor performance. GCMC performs better than NeuMF, highlighting the benefits of leveraging graph structure. However, GCMC underperforms LightGCN because it not explicitly leverage high-order connectivities between users and items. This showcases the strength of GCN and the importance of incorporating high-order information into representation learning. Moreover, DGCF performs better than LightGCN by leveraging the disentangled representation technique to model users’ diverse intents. Both IMP-GCN and NCL consistently outperform all the above GCN-based methods over all the datasets. This is mainly due to their ability to filter out noisy information from neighboring nodes effectively. In particular, IMP-GCN’s performance superiority over LightGCN and DGCF demonstrates the importance of distinguishing nodes in high-order neighbors in the graph convolution operation.
The second block is the GCN-based methods utilizing user reviews as side information. MMGCN is the first method combining side information with GCN, it exploits the high-order information of various modalities. However, it falls short of GRCN in terms of effectiveness. The reason is that the GRCN could adaptively adjusting the false-positive edges of the interaction graph to discharge the ability of GCN. LATTICE outperforms GRCN on Baby and Clothing by constructing item-item graphs based on item similarity in the side information. BM3 consistently performs better than all general CF baselines over Baby and Clothing. It should be credited to the joint optimization of multiple objectives to learn the representations of users and items.
The third block illustrates that review-based methods primarily use user textual information as a key semantic feature. DeepCoNN, which employs Convolutional Neural Networks, designs two coupled networks to learn user and item representations from reviews, outperforming NeuMF due to its use of semantic information. NARRE enhances recommendations by using attention mechanisms to weigh the significance of different reviews, showing particular improvement on the sparse Clothing dataset. RGCL further incorporates review data by building a review-aware user-item graph, outperforming all three review-based methods. To validate the ability of LLMs in understanding reviews, we designed LightGCNLDA, which combines LDA with LightGCN. Similar to our approach, it uses topic-level information to construct user-item graphs and then learns user and item representations through GCN techniques. As a result, it outperforms all review-based models across all datasets.
The fourth block consists of two LLM-based methods: LlamaRec and RLMRec. LlamaRec is a two-stage framework that first retrieves candidate items and then ranks them using an LLM. It outperforms NeuMF due to the advanced capabilities of LLMs in understanding complex patterns and relationships within the data. RLMRec, on the other hand, outperforms LlamaRec by leveraging a more comprehensive data approach and generating richer user and item profiles. While LlamaRec focuses solely on user interaction data, RLMRec combines interaction data with textual information such as product descriptions, user reviews, and product titles. This enables RLMRec to create detailed, contextually enriched profiles, offering a deeper understanding of user preferences and item characteristics.
SAGCN consistently provides the best performance across all datasets. In particular, the improvements of SAGCN over the strongest competitors $w.r.t$ NDCG@10 are 15.59%, 15.89%, 21.77%, and 8.50% in Office, Baby, Clothing, and Goodreads, respectively. We credit this to the combined effects of the following three aspects. Firstly, we accurately extract structured semantic aspect-aware reviews using LLM to understand user behaviors at a finer level, which also avoids the noise in raw reviews. Our proposed method outperforms LightGCNLDA, highlighting the superior understanding capabilities of LLMs. Additionally, the semantic aspects generated by LLMs offer better interpretability compared to the latent topics extracted by LDA. Secondly, the SAGCN performs graph convolution on various semantic-aspect-based graphs, which can alleviate the over-smoothing problem (Liu et al., 2021) and avoid negative information from unreliable neighbor nodes (Wei et al., 2020). Thirdly, we extract fine-grained user behaviors for each semantic aspect to bridge the modality gap between interaction and textual features. Different with existing methods that integrate textual features with user-item interaction data for representation learning, our approach constructs structured user-item interactions from user reviews via LLM, which are associated with various semantic aspects. These structured interactions are then leveraged for the representation learning of users and items.
4.3. Performance Comparison with Pure LLM
To further validate the effectiveness of our approach, we implemented a method named LLM, built upon LLM for comparative analysis. However, due to the high resource and time consumption of LLM, we evaluated the method by randomly selecting 200 users from each dataset. To construct the testing data, we randomly selected one positive item from the history items of the original testing data. We then sampled 9 negative items not interacted with by the user as the testing set for each user. Next, we converted the historical data and items in the testing set for each user into textual format using a prompt template. Finally, we fed the converted text into LLM to generate scores for the items in the testing set. These items are then ranked according to the scores generated by the LLM. Our prompt template is as follows:
I want you to rate every candidate product’s historical record of purchased habits. You are encouraged to learn his preferences from the historical records he has purchased. Here are The historical interactions of a user include: { history }. Now, how will the user rate these candidate products? (1 being lowest and 5 being highest) { candidates }. Importantly, the interacted items should have been excluded from the rating. Finally, Only output rating item list, which template is: 1. Swingline GBC UltraClear Thermal Laminating Pouches, Menu Size, 3 Mil, 25 Pack (item id: B00006IA2K) - Rating: 4.0 stars
Where {history} consists of item title and user rating score; {candidates} is composed of item titles and item id. Besides, we also use the Random method as a comparison method, which randomly ranks the items in the testing set.
As illustrated in Tables 3 and 4, the performance of the Random method is inferior as it randomly recommends the items to users. The LLM method achieve better performance than Random. This is because it has the ability to recognize patterns and correlations in data. This capability allows it to predict user preferences and make more informed recommendations based on observed patterns in user behavior or item characteristics. However, its performance does not significantly exceed that of Random, especially in terms of Recall@1 and NDCG@1. This can be attributed to the LLM not having been trained or fine-tuned for the specific recommendation task. Experimental results demonstrate the effectiveness of our proposed approach, which integrates LLM with conventional recommendation techniques. A remarkable advantage of SAGCN is its aspect-aware graph, which allows for a finer understanding of complex interactions between users and items, further augmented by the sentiment analytical power of LLM.
Table 3. Performance Comparison of Random, LLM* and our method on Office.
Datasets Office
Metrics k=1 k=3 k=5 k=7
Recall NDCG Recall NDCG Recall NDCG Recall NDCG
Random 0.0835 0.0835 0.2926 0.2012 0.5261 0.2962 0.7203 0.3632
LLM 0.1086 0.1086 0.5803 0.3903 0.7660 0.4657 0.9122 0.5163
SAGCN 0.4880 0.4880 0.7470 0.6371 0.8763 0.6838 0.9669 0.7289
Table 4. Performance Comparison of Random, LLM* and our method on Clothing.
Datasets Clothing
Metrics k=1 k=3 k=5 k=7
Recall NDCG Recall NDCG Recall NDCG Recall NDCG
Random 0.0841 0.0841 0.2848 0.2208 0.4474 0.2867 0.6186 0.3458
LLM 0.0912 0.0912 0.4860 0.2997 0.7692 0.4170 0.8913 0.4669
SAGCN 0.4823 0.4823 0.7239 0.6176 0.8273 0.6603 0.9245 0.6908
4.4. Effect of Layer Numbers
(a) Recall on Office
(b) Recall on Clothing
(c) Recall on Baby
(d) Recall on Gooodreads
(e) NDCG on Office
(f) NDCG on Clothing
(g) NDCG on Baby
(h) NDCG on Gooodreads
Figure 5. Performance Comparison between SAGCN and competitors at different layers on Office, Clothing, Baby, and Goodreads. Notice that the values are reported by percentage with ’%’ omitted.
4.5. Handling of Noisy Neighboring Nodes
We conducted a case study to illustrate how our approach handles noisy neighboring nodes. As shown in Fig. 6, we selected user $u_{0}$ as the target node and chose some of its first, second, and third-order neighboring nodes to construct a graph displaying their connections. In the original graph, when performing graph convolutions, information from all neighboring nodes is passed to the target node $u_{0}$ through the same graph structure. In contrast, our approach, which leverages LLMs to analyze the semantic aspects mentioned in each review, constructs aspect-aware graphs based on these results. This means that certain neighboring nodes may only appear in specific aspect-aware graphs. For instance, in user $u_{2546}$’s review of $i_{1534}$, only the aspect of Price was mentioned, while Ease-of-Use was not. As a result, when constructing the graph, node $i_{1534}$ only appears in the Price aspect-aware graph and passes information to the target node through this specific graph structure.
By constructing aspect-aware graphs based on specific semantic aspects identified in user reviews, our method ensures that only the most relevant information is passed from neighboring nodes to the target node. In traditional graph convolution methods, all neighboring nodes contribute information equally, even if their relevance to the target node is minimal or unrelated, which introduces noise and dilutes the signal from more pertinent nodes. With our approach, neighboring nodes that are not relevant to a specific aspect (e.g., Ease-of-Use or Price) are excluded from the corresponding aspect-aware graph. This selective filtering reduces the contribution of irrelevant or less important nodes, allowing the target node to receive more focused and meaningful information. As a result, the quality of the information passed to the target node improves, leading to more accurate representations and better overall model performance.
Figure 6. Illustration of Aspect-Aware Graph Construction for Reducing Noisy Neighboring Nodes in Graph Convolutions. User $u_{0}$ is the target node, and we selected some of its first, second, and third-order neighboring nodes to show their connectives in the original graph and aspect-aware graphs.
4.6. Effect of Embedding Dimensions
(a) Recall on Office
(b) NDCG on Office
(c) Recall on Clothing
(d) NDCG on Clothing
Figure 7. Impact of embedding dimensions on the performance of our proposed method, SAGCN, in comparison with three GCN-based methods (LightGCN, DGCF and RGCL). Notice that the values are reported by percentage with ’%’ omitted.
In this section, we analyze the impact of embedding dimensions on the performance of our proposed method, SAGCN, in comparison with three GCN-based methods (LightGCN, DGCF and DCCF). We experiment with embedding dimensions ranging from 64, 128, 256, 384, 512, evaluating how the size of the embeddings influences model accuracy and overall performance.
From the results shown in Fig. 7, we observe a general improvement in performance across all methods as the embedding dimensions increase. However, for LightGCN, DGCF, and RGCL, performance tends to plateau after reaching 128 or 256 dimensions. In contrast, our method, SAGCN, initially performs poorly at 64 dimensions but shows rapid performance growth as the dimensions increase, particularly beyond 384 dimensions. This improvement can be attributed to the aspect-aware graphs constructed using LLMs to analyze user reviews. Our approach transforms simple user interactions into multiple interaction patterns, significantly enriching the informational content of the interaction data, especially after constructing the aspect-aware graphs. As a result, a larger feature space is required to effectively learn user and item representations from these enriched interactions. The increased information not only enhances the performance of our method but also necessitates a larger feature space to capture the complexity of these interactions.
(a) Recall on Office
(b) NDCG on Clothing
(c) Recall on Baby
(d) Recall on Goodreads
(e) NDCG on Office
(f) NDCG on Clothing
(g) NDCG on Baby
(h) NDCG on Goodreads
4.7. Semantic Aspects Analysis
@@
Table 5. The extracted semantic-aware aspects and the number of their associated interactions over four datasets.
@@
Goodreads Satisfaction Relevance Appearance Support Price Usefulness Quality Creativity
34,774 31,686 19,707 19,366 14,903 14,111 11,394 10,254
@@ 4.7.1. Semantic Aspect-Aware Interactions Analysis
In our experiment, we selected eight aspects per dataset by merging similar aspects and eliminating those with very low occurrence. Table 5 reported the number of interactions for different semantic aspects across four datasets.
Based on our observations of various recommendation scenarios, it is evident that users have distinct concerns regarding different aspects. For instance, in reviews of office products, convenience is a common concern, while fit is only mentioned in clothing reviews. To rank the semantic aspects in each recommendation scenario, we analyzed their interaction numbers. This helps us identify that the importance of different semantic aspects varies depending on the number of interactions. For example, in reviews of Office products, Durability was mentioned more frequently than Design. On the other hand, in reviews of Baby, Design was mentioned more frequently than Durability. Note that the most frequently mentioned aspect being Quality in the Office, Baby, and Clothing datasets. For these types of products, customers often prioritize Quality because it directly influences their decision-making. In contrast, for the GoodReads, the most mentioned aspect is Satisfaction, which aligns with the subjective nature of book reviews. Readers tend to express overall satisfaction with the reading experience, which encompasses personal enjoyment, engagement with the story, or alignment with expectations.
@@ 4.7.2. Effect of Aspect Numbers
To investigate the effect of semantic aspect numbers, we conducted experiments across all four datasets. Note that all aspects were ranked according to the number of interactions, we selected a number of aspects based on their ranking. Our findings, shown in Fig. 8, suggest that increasing the semantic aspect numbers can improve the performance of SAGCN. In other words, SAGCN achieves its peak performance when leveraging all aspects, indicating that it can benefit from various aspects. Furthermore, we found that SAGCN with eight aspects showed a greater improvement in performance than it with six aspects. This could be due to the fact that the added two aspects might provide more supplementary information to the other six. When combined, the full set of features gives the model a richer understanding of user behaviors.
@@ 4.8. Interpretability Analysis
(c) Recall on Baby
(d) Recall on Goodreads
(e) NDCG on Office
(f) NDCG on Clothing
(g) NDCG on Baby
(h) NDCG on Goodreads
@@
Figure 9. Performance comparison of SAGCN on different semantic aspects across all four datasets. Notice that the values are reported by percentage with ’%’ omitted.
(a) User (u3547)
(b) Item (i2210)
(c) User (u3547)
@@
Figure 10. Three examples of preference scores across different aspects in Office. a) Preferences of user $u3547$ for item $i836$ and $i1322$ on different semantic aspects. b) Preferences of user $u4842$ and $u2071$ for item $i2210$ on different semantic aspects. c) Preferences of user $u3547$ for item $i836$ and $i1322$ on different latent factors. Notice that the values are reported by percentage with ’%’ omitted.
@@
In this section, we examined the interpretability of our model from two different perspectives. First, our model can interpret the contribution of each semantic aspect in a recommendation scenario. To examine the contribution of different aspects, we predicted the users’ preferences using individual aspect embeddings of the users and items. Fig. 9 shows the following interesting observations: (1) Different semantic aspect features have varying contributions in recommendation. For instance, Quality has a greater impact than Design for Office. (2) The number of semantic-aware interactions is not the main contributory factor. Surprisingly, even though the number of interactions for Price is the least, it still has the most contribution to our model. This could be because Price introduces more valuable interactions than the other aspects.
@@
We can further provide a detailed explanation by providing the scores of various semantic aspects for each user-item pair. In Fig. 10 (a), we computed the preference score of user $u3547$ for items $i836$ and $i1322$ from different semantic aspects in Office. From the result, we can conclude that the user has similar but different preferences on different aspects of different items. Moreover, the user would be interested in these two items mainly due to their Convenience and Quality. Additionally, the aspect of Comfort in $i1322$ might be more appealing to the user than $i836$. Similarly, we computed the preference score of users $u4842$ and $u2071$ for item $i836$, which is illustrated in Figure 10 (b). These two users have diverse preferences for the same item. Specifically, user $u4842$ focuses on Design, Price and Convenience. In contrast, user $u2071$ pays more attention to the Design, Price, and Durability. These experimental results also demonstrate the interpretability of our proposed approach.
Overall, we further emphasize our contribution to interpretability compared to traditional methods using disentanglement techniques. As shown in the fig. 10 (c), in disentanglement-based methods, each representation chunk is associated with a latent factor that cannot be understood by humans. In other words, while we can identify which factor contributes more to user preferences, we have no understanding of what that factor actually represents. In contrast, our approach associates the representations learned from different graphs with specific semantic aspects, which are directly understandable by humans. This demonstrates the advantage of our method in enhancing interpretability, as it provides clear, human-readable explanations for the factors driving user preferences.
4.9. Feature Independence Analysis
(a) Similarity scores among different aspects (Office)
(b) Similarity scores among different aspects (Clothing)
Figure 11. Visualization of the feature independence of semantic aspects.
We present a visualization of the feature independence of different semantic aspects in Office and Clothing, as shown in Fig. 10. To produce this, we first sampled one user from the user set and then calculated the cosine similarity scores among different aspects based on the user’s representation vector. The heatmap in Fig. 10 shows that the similarity values are close to zero between different aspects. This implies that the representations associated with different semantic aspects are independent. This demonstrates that the SAGCN model can learn robust representations with the semantic aspects extracted via LLM, enabling more accurate item recommendations for users. An additional noteworthy finding is that the representations are unaffected by the potential bias stemming from the number of semantic aspect-aware interactions. For example, as indicated in Tab. 5, there is an imbalance in the number of aspect-aware interactions(e.g., the Quality aspect-aware interactions occupy a large proportion in all aspect-aware interactions on both Office and Clothing). In previous data-driven recommendation models, such an imbalance might lead the model to overly focus on these dominant aspects. However, the SAGCN approach leverages the aspect-aware interactions solely for constructing the semantic-aware graph, meaning that the quantity of these interactions does not sway the aspect-associated representations. This further demonstrates that our proposed model can obtain more robust and better representations by capturing the fine-grain relations between each user and item at the semantic aspect level.
5. Conclusion and Future Work